TK50 Tape Drive

Various Random Folks October 27, 1992

How to use the TK50 Tape Drive on yaz-pistachio:

1. First blow any dust out of the tape drive...I’m not kidding....

2. Get a nice TK50 tape (not a TK70).

3. If you want to write to the tape, make sure the little orange spot on the tape cartridge isn’t showing; if it shows, the tape is write-protected.

4. Place tape in drive. Close door (flap). WAIT.

5. Press red button. WAIT. The red button takes the tape on/offline.

6. Invoke appropriate patron saint, diety, or guardian angel. WAIT.

7. Listen for grinding sounds to stop. WAIT.

8. Use “tar tv”, or use “ansitape -tv”, to list the files on the tape. If you don’t get errors here, you should be able to read the tape.

9. Use “tar xv [filenames]”, or use “ansitape -xv [filenames]”, to extract filenames from the tape. See man pages for “tar” and “ansitape” for other file transfer operations.

10. All tape accesses will take some time - the drive is SLOW. WAIT.

11. The green light signifies that the tape is being accessed. The green light should be off before you can remove a tape from the drive. If the red light is blinking slowly and the green light is blinking quickly, the tape is rewinding. WAIT.

12. Before removing tape cartridge from the drive, press red button. WAIT.

13. Open door (flap). Remove tape, hopefully.

14. Praise appropriate patron saint, diety, or guardian angel invoked in step 6 if successful.

15. Call Anne Salemme at x3-0129 if you have problems.